OpenStudy (anonymous):

simplify: (20x^3)1/2 + (45x^3)1/2

OpenStudy (anonymous):

\[\frac{ 20x^3 }{ 2 } + \frac{ 45x^3 }{ 2 } = \frac{ 20x^3 + 45x^3 }{ 2 } = \frac{ 65x^3 }{ 2 }\]

OpenStudy (anonymous):

sorry the right form is : (20x^3)^1/2 + (45x^3)^1/2) sorry these symbols get confusing

OpenStudy (anonymous):

\[(20x^3)^{1/2} + (45x^3)^{1/2} = \sqrt{20x^3} + \sqrt{45x^3} = \sqrt{20}\sqrt{x^3} + \sqrt{45}\sqrt{x^3}\]\[= \sqrt{4}\sqrt{5}x^{3/2} + \sqrt{9}\sqrt{5} x^{3/2} = 2x^{3/2}\sqrt{5} + 3x^{3/2}\sqrt{5} = x^{3/2}\sqrt{5} (2 + 3)\]\[= 5x^{3/2}\sqrt{5}\]